# Tree Attention: Topology-Aware Decoding for Long-Context Attention on GPU Clusters

### v2

@@ 1.  Introduction
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu and Dao (2023); Dao and Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024). Other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory-efficient attention Rabe and Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works, Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. This is the regime of primary interest of this paper. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices, instead of linearly in alternatives such as Ring Attention Liu et al. (2023). Our topology aware approach illustrated in Fig. 1. significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
@@
  • We provide a mathematical form for the energy function of self attention.
  • From this theory, we develop an algorithm for parallelizing the attention computation across devices, leveraging tree reduction topology.
  • Finally, we show empirically that our algorithm provides asymptotic speedups when decoding across many GPUs on long sequence lengths as well as large reductions in communication volume and peak memory usage.
@@ 2.  Related Work
To advance beyond this, our energy function depends on an auxiliary source vector $\zeta$ (see section 4.). Taking the gradient with respect to $\zeta$ yields the exact self-attention operation, as we show below. Consequently, this formulation enables us to express self attention with no weight tying.
Another notable related work is Feng et al. (2024) where the authors also made similar observations as we do in section 5. about how the associative operations within the attention computation can be efficiently parallelized to motivate an attention based modified RNN architecture for sequence modeling.
@@ 3.  Self-Attention
The self-attention operation can be represented as a set of dot product similarity searches between queries and keys. These similarity scores are then reduced along the sequence axis and softmaxed, so that for a given query, there is a probability distribution of the similarities of each given key. We then take the expectation of the value vectors against this distribution. We denote the queries assigned to a sequence of length $N$ as $\{q_{a},a=1,\cdots N\}$, where each query is a vector of size $d$ that stands for hidden dimension: $q_{a}\in\mathbb{R}^{d}$ and similarly the keys and values $\{(k_{a},v_{a}),a=1,\cdots N\}$. Attention can be written as
$z^{a}=\sum^{N}_{i=1}\text{softmax}(q_{a}\cdot k_{i}^{T})v_{i}$
@@ 4.  Self-Attention as a Gradient
Here, we show how the self-attention operation can be written as the gradient of a certain scalar function. In particular, we define a scalar function that depends on the keys, queries, values and additionally, on an auxiliary vector that we refer to as the source $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self-attention operation. We need the source in order to write down the generating function of the moments of the distribution above. The source is also the variable with respect to which we can Taylor-expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
@@ 4.1.  The energy function
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We proceed focusing on the single head case, as it makes the presentation simpler, and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix 11.
@@ 4.2.  Bayesian interpretation
Moreover, such a procedure enables us to identify the energy based model associated with the self-attention function.
@@ 5.1.  Computing the energy efficiently
We can prove that this associative property allows these reduction to be performed efficiently in parallel with logarithmic time complexity, provided we have adequately many parallel workers:
@@ 5.2.  Efficient parallel decoding
Notice here that by storing $O,m$ for backward, the onlt remaining reduction operation that needs to be performed is the one in line 5 of the above algorithm. This single reduction takes $O(N/p)$ time to compute the local sums on each device and $\log p$ time to communicate and combine partial results therefore we get the same asymptotic complexity as the logsumexp calculation.
In practice we implement the forward and gradient w.r.t. $\zeta$ in a single function which returns both the value and the gradient of the energy function. We can therefore put together Algorithms 1 and 2 into the following efficient parallel decoding algorithm 3:
@@
4:Obtain the local numerator and denominator by then computing: $\textrm{n}=\textrm{o}*\exp(\textrm{lse}),\textrm{d}=\exp(\textrm{lse})$.
5:Recompute the global max $m=\textrm{{Allreduce}}(\textrm{max},\textrm{d})$.
6:Compute global numerator and denominator $\textrm{n}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{n}e^{-m}),\textrm{d}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{d}e^{-m})$.
@@ 5.3.  Implementation and topology-awareness
Figure 2: NCCL Send/Recv between two H100 GPUs intra-node and inter-node. GPU clusters offer a two-tier topology whereby intra-node bandwidth is significantly higher than inter-node. Algorithms such as Tree Attention exploit this topology by reducing internode communication requirements, enabling better overlap of communication with computation.
@@ 6.1.  Latency
To empirically test the theoretical benefits of our Tree Attention method, we compute latency by measuring the time required to perform decoding for different sequence lengths and varying number of H100 nodes. We compare Tree Attention to our own Ring Attention execution times in Fig. 4(a). Both methods use FlashAttention-2 Dao (2023) for the individual-GPU attention computation. For our experiments, we benchmark on a standard attention block consisting of 16 heads of dimension 128 across different sequence lengths.
@@ 7.  Discussion and Conclusion
Our introduction of a unique energy function for self attention develops interesting connections between attention and other related models such as Hopfield networks and the general notion of associative memories. Given this energy function, it is possible to mathematically analyze its loss landscape and dynamics, as well as begin to understand how the attention operation could potentially be improved. Additionally, since the energy function can be framed as a variational free energy, it becomes possible to interpret the attention operation through a Bayesian lens and understand the generative model that is implicit in its operation. We hope that these mathematical derivations aid research in understanding the properties of transformer networks, as a deep understanding of the Hopfield energy function has done for Hopfield networks.

### v3

@@ 1.  Introduction
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu and Dao (2023); Dao and Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024). Other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory-efficient attention Rabe and Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. This is the regime of primary interest of this paper. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices, instead of linearly in alternatives such as Ring Attention Liu et al. (2023). Our topology-aware approach illustrated in Fig. 1. significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
@@
  • We provide a mathematical form for the energy function of self-attention.
  • From this theory, we develop an algorithm for parallelizing the attention computation across devices, leveraging tree-reduction topology.
  • Finally, we show empirically that our algorithm provides asymptotic speedups when decoding across many GPUs on long sequence lengths, as well as large reductions in communication volume and peak memory usage.
@@ 2.  Related Work
To advance beyond this, our energy function depends on an auxiliary source vector $\zeta$ (see section 4.). Taking the gradient with respect to $\zeta$ yields the exact self-attention operation, as we show below. Consequently, this formulation enables us to express self-attention with no weight tying.
Another notable related work is Feng et al. (2024), where the authors made similar observations as we do in section 5. about how the associative operations within the attention computation can be efficiently parallelized to motivate an attention-based modified RNN architecture for sequence modeling.
@@ 3.  Self-Attention
The self-attention operation can be represented as a set of dot product similarity searches between queries and keys. These similarity scores are then reduced along the sequence axis and softmaxed, so that for a given query, there is a probability distribution of the similarities of each given key. We then take the expectation of the value vectors against this distribution. We denote the queries assigned to a sequence of length $N$ as $\{q_{a},a=1,\cdots N\}$, where each query is a vector of size $d$ that stands for hidden dimension, $q_{a}\in\mathbb{R}^{d}$, and similarly the keys and values $\{(k_{a},v_{a}),a=1,\cdots N\}$. Attention can be written as
$z^{a}=\sum^{N}_{i=1}\text{softmax}(q_{a}\cdot k_{i}^{T})v_{i}\,.$
@@ 4.  Self-Attention as a Gradient
Here, we show how the self-attention operation can be written as the gradient of a certain scalar function. In particular, we define a scalar function that depends on the keys, queries, values and additionally on an auxiliary vector that we refer to as the source $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self-attention operation. We need the source in order to write down the generating function of the moments of the distribution above. It is also the variable with respect to which we can Taylor-expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
@@ 4.1.  The energy function
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We proceed focusing on the single-head case, as it makes the presentation simpler, and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix 11.
@@ 4.2.  Bayesian interpretation
Moreover, such a procedure enables us to identify the energy-based model associated with the self-attention function.
@@ 5.1.  Computing the energy efficiently
We can prove that this associative property allows these reductions to be performed efficiently in parallel with logarithmic time complexity, provided we have adequately many parallel workers:
@@ 5.2.  Efficient parallel decoding
Notice here that by storing $O,m$ for backward, the only remaining reduction operation that needs to be performed is the one in line 5 of the above algorithm. This single reduction takes $O(N/p)$ time to compute the local sums on each device and $\log p$ time to communicate and combine partial results, and therefore we get the same asymptotic complexity as the logsumexp calculation.
In practice, we implement the forward and gradient w.r.t. $\zeta$ in a single function which returns both the value and the gradient of the energy function. We can therefore put together Algorithms 1 and 2 into the following efficient parallel decoding Algorithm 3:
@@
4:Recompute the global max $m=\textrm{{Allreduce}}(\textrm{max},\textrm{lse})$.
5:Obtain the local numerator and denominator by then computing: $\textrm{n}=\textrm{o}*\exp(\textrm{lse}-m),\textrm{d}=\exp(\textrm{lse}-m)$.
6:Compute global numerator and denominator $\textrm{n}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{n}),\textrm{d}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{d})$.
@@ 5.3.  Implementation and topology-awareness
Figure 2: NCCL Send/Recv between two H100 GPUs intra-node and inter-node. GPU clusters offer a two-tier topology whereby intra-node bandwidth is significantly higher than inter-node. Algorithms such as Tree Attention exploit this topology by reducing inter-node communication requirements, enabling better overlap of communication with computation.
@@ 6.1.  Latency
To empirically test the theoretical benefits of our Tree Attention method, we compute latency by measuring the time required to perform decoding for different sequence lengths and varying number of H100 nodes. We compare Tree Attention to our own Ring Attention execution times in Fig. 4(a). Both methods use Flash Attention 2 Dao (2023) for the individual-GPU attention computation. For our experiments, we benchmark on a standard attention block consisting of 16 heads of dimension 128 across different sequence lengths.
@@ 7.  Discussion and Conclusion
Our introduction of a unique energy function for self-attention develops interesting connections between attention and other related models such as Hopfield networks and the general notion of associative memories. Given this energy function, it is possible to mathematically analyze its loss landscape and dynamics, as well as begin to understand how the attention operation could potentially be improved. Additionally, since the energy function can be framed as a variational free energy, it becomes possible to interpret the attention operation through a Bayesian lens and understand the generative model that is implicit in its operation. We hope that these mathematical derivations aid research in understanding the properties of transformer networks, as a deep understanding of the Hopfield energy function has done for Hopfield networks.